OpenStudy (anonymous):

when heated, KCIO3 decomposes into KCI and o2. If this reaction produced 14.5 grams of KCI, how much O2 was produced (in grams)?

OpenStudy (anonymous):

ok so first write ballanced chemical equation and then we will go on...

OpenStudy (anonymous):

2 KClO3 -> 2 KCl + 3 O2 so you dont know basics... lets move further on: we write down what we have and what we want: m(KCl) = 14,5 g --------------- m(O2) = ? now when you learn this routine in solving this kind of problems you wont have any more problems just follow these steps: n(O2)/n(KCl) = 3/2 n(O2) = 3/2 n(KCl) n(O2) = 3/2 * m(KCl)/M(KCl) now since n=m/M m(O2) = 3/2 * [m(KCl)/M(KCl)] * M(O2) any questions?

OpenStudy (moonlight93):

can someone explain more. I still don't get it.